# Navigating Distribution Shifts in Medical Image Analysis: A Survey

### v1

@@ IV-B Semi-supervised Joint Training
IV-B1 Data Management
Pseudo-labeling. Pseudo-labeling is a powerful technique in semi-supervised joint training that leverages large volumes of unlabeled data to enhance model training. This method involves generating artificial labels for unlabeled data based on the most confident predictions of the model, thereby expanding the training dataset effectively. Specifically, it employs the learned features from training data to generate pseudo-labels for unlabeled data. In domain adaptation, its effectiveness hinges on prioritizing model’s predictions exceeding a defined threshold that are more likely to be accurate:
$\hat{y}_{u}=\begin{cases}\arg\max q_{\theta}(x_{u}),&\text{if }\max q_{\theta}(x_{u})>\tau\\ \text{ignore},&\text{otherwise}\end{cases}$ (1)
where $x_{u}$ represents an unlabeled sample from the target domain, $q_{\theta}(x_{u})$ denotes the model’s probabilistic predictions, $\hat{y}_{u}$ is the pseudo-label assigned to $x_{u}$, and $\tau$ is a threshold defining the confidence level above which the labels are considered reliable. The pseudo-label loss is computed as:
$\mathcal{L}_{u}=\sum_{x_{u}\in\mathcal{X}_{u}}\mathbb{I}\left(\max q_{\theta}(x_{u})>\tau\right)\cdot\mathcal{L}(\hat{y}_{u},q_{\theta}(x_{u}))$ (2)
@@ IV-B1 Data Management
$\mathcal{L}_{total}=\mathcal{L}_{l}(\mathcal{Y}_{l},q_{\theta}(\mathcal{X}_{l}))+\beta\mathcal{L}_{u}(\mathcal{\hat{Y}}_{u},q_{\theta}(\mathcal{X}_{u}))$ (3)
@@
To mitigate the risk of error propagation, which can occur if incorrect labels are used for training, enhancements are made to ensure the quality of these pseudo-labels. For example, [43] employs transformation-invariant, highly-confident predictions in the target dataset for self-training purposes, ensuring that the model is less likely to learn from noisy, less reliable labels. Meanwhile, [44] enhances the robustness of pseudo-labeling by calculating the variance between the original image and its Fourier-transformed counterpart, providing a more stable basis for generating reliable pseudo-labels. These strategies significantly improve the utility of pseudo-labeling, making it a vital tool for utilizing unlabeled data in domain adaptation.
IV-B2 Model Design
Self-ensembling. Self-ensembling is an advanced learning strategy that effectively exploits both labeled and unlabeled data with the consistency between models. This method trains multiple versions of a model, each subjected to distinct input perturbations, and employs consistency regularization to ensure uniform predictions across these variations. Typically, this technique under SSDA setting is implemented via a “teacher-student” model, where a stable, pre-trained “teacher” model $q_{\theta}^{tc}$ guides a less-trained “student” model $q_{\theta}^{st}$. The overall training process is governed by two primary loss functions:
$\mathcal{L}_{{total}}$ $=\mathcal{L}_{l}(\mathcal{Y}_{l},q_{\theta}^{tc}(\mathcal{X}_{l}))$
$+\lambda\mathcal{L}_{{cons}}(q_{\theta}^{tc}(\mathcal{X}_{u})),q_{\theta}^{st}(\mathcal{X}_{u}^{aug}))$ (4)
@@ IV-B2 Model Design
Following this framework, [45, 46] introduce this techniques into achieveing semi-supervised domain adaptation. The core innovation of these frameworks lies in the strategic employment of dual-teacher models: one teacher model enhances intra-domain knowledge through self-ensembling techniques, while the other facilitates inter-domain knowledge transfer using image translation models such as CycleGAN [47]. This approach leverages the consistency of model outputs across different views of the same data, enhancing the model’s ability to generalize across diverse scenarios.
Adversarial Learning. Adversarial leaning is fundamentally embodied by Generative Adversarial Networks (GANs). It can be conceptualized as a game between two players:
$\min_{G}\max_{D}V(D,G)$ $=\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]$
$+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (5)
where generator (G) aims to produce data that is indistinguishable from real data by transforming input noise $z$, sampled from a noise distribution $p_{z}(z)$, and discriminator (D) aims to correctly classify real data $x$ and generated data $G(z)$. Real data $x$ is sampled from the true data distribution $p_{\text{data}}(x)$. This interaction forms a min-max game where the generator seeks to deceive the discriminator into accepting its outputs as real, while the discriminator improves at identifying the differences between real and generated data. Through this adversarial process, $G$ refines its outputs to reduce discrepancies, indirectly generating domain-invariant features for domain adaptation. As one notable method, COVID-DA [48] is designed to distinguish between closely related conditions such as pneumonia and COVID-19, particularly when labeled data is scarce. This method uses a unique classifier separation scheme along with an adversarial network to overcome the task difference and domain discrepancy simultaneously.
Novel Training Strategies. [49] explores the richness of multi-modal data through a novel asymmetric co-training approach. By segmenting the learning process into two distinct components that each addresses specific aspects of domain adaptation and semi-supervised learning task, this strategy avoids the domination of the source domain data thus facilitates more effective domain adaptation.
IV-B3 Optimization Strategy
Metric Learning. Metric Learning within the context of semi-supervised joint training is distinctly innovative. In this scenario, [50] adopts a metric learning strategy characterized by a disentangled paradigm. This approach separates style and content into distinct embedding spaces. Such separation facilitates independent contrastive learning for each aspect, allowing the model to adapt more effectively to variations in data distributions.
IV-C Unsupervised Joint Training
@@
IV-C1 Data Management
Cross-Modal Translation. Cross-modal translation, employing techniques such as Generative Adversarial Networks (GANs) and frequency-based methods, is pivotal in transforming how we address domain differences by converting data from the source domain $S$ to closely resemble the target domain $T$. This transformation is formalized as:
$\mathcal{\tilde{X}}=G(\mathcal{X}_{s};\theta_{G})\quad\text{where}\quad G:\mathcal{D}_{s}\rightarrow\mathcal{D}_{t}$ (6)
@@ IV-C1 Data Management
$\min_{\theta_{G}}d(\mathcal{\tilde{X}},\mathcal{X}_{t})=\min_{\theta_{G}}d(G(\mathcal{X}_{s};\theta_{G}),\mathcal{X}_{t})$ (7)
@@
The application of GANs has evolved significantly, beginning with methods like the pixel-to-pixel (pix2pix) translation and advancing to more complex implementations. For instance, [51] utilizes an end-to-end unsupervised method for enhancing contrast in cataract fundus image based on pix2pix framework. Then, MADGAN [52] breaks the constraints of paired images, contributing to anomaly detection in complex brain structures. SASAN [53] takes a further step that incorporating self-attention modules in its GANs, enhancing focus on specific anatomical details during image translation. Subsequently, the utilization of CycleGANs [47] marks another significant advancement, enabling unpaired image translations for cross-domain chest X-ray disease recognition [54, 55] and hip joint bone segmentation [56]. This translation process is further refined in [57] with dual-scheme (source-target/target-source) fusion and [58] with attention mechanism. Integrating disentangled representations into GAN frameworks, as seen in [59, 60, 61, 62, 63, 64], significantly advances domain adaptation by separating content from style, enhancing adaptation efficiency. Complementing the adversarial nature of GANs, frequency-based methods [65, 66] introduce a novel perspective. They assume that the style information is stored in low frequency components and high frequency components represents more structural information, and thus translate the images by replacing the low frequency components. Finally, techniques like singular value decomposition for noise adaptation in retinal OCT images [67] highlight the innovation and adaptability in this field, which is tailored to specific imaging modalities or diagnostic requirements.
@@
Pseudo-labeling. In unsupervised joint training, as all target data labels are unknown, it becomes more challenging to make accurate pseudo-label predictions using traditional techniques. Research has since advanced the pseudo-labeling concept by integrating pseudo-labeling and adversarial learning to enhance the process [68]. Subsequent studies have built on this foundation, each offering unique improvements to address issues such as noisy labels [69] and enhancing label reliability through methods like iterative self-training [70], contrastive learning [71], and entropy constraints [72]. The specialized applications of pseudo-labeling are further explored in studies [73, 74, 75]. For example, [73] focuses on nuclei instance segmentation and classification, utilizing pseudo-labels derived from prototype features. [74] breaks new ground in cell detection with a pseudo-cell-position heatmaps. [75] innovates by incorporating pseudo-labeling into tagged-to-cine MRI synthesis task, employing a Bayesian uncertainty mask for selective pseudo-label generation.
IV-C2 Model Design
@@
Adversarial Learning. Adversarial Learning is widely used for the implicit alignment between domains at feature or/and pixel level due to the absence of target labels. At the feature level, techniques such as the plug-and-play adversarial domain adaptation network (PnP-AdaNet) [76] aligns features across different scales for segmentation tasks. Similarly, [77] aligns extracted contents for cross-modality segmentation. Other studies focus on prediction space alignment at the pixel level for various medical imaging tasks [78, 79, 80, 81]. Integrated approaches that apply adversarial training at both feature and output levels are explored in studies like [24, 82, 83, 84, 85, 86, 87]. Innovations in this field also include enhanced discriminators and local discriminators that focus on specific region alignment, introducing spatial-aware and class-specific attentions to refine the adversarial loss and improve model adaptability across domains [88, 89, 90, 91].
@@
Self-ensembling. Initial studies by [92] applied self-ensembling to gray matter MRI segmentation. Subsequent applications include breast MRI segmentation [93] and pose estimation in operating rooms [94]. More advanced techniques combine adversarial training and self-ensembling for addressing domain shifts in cross-institutional gliomas studies [95], optic disc and cup segmentation [96] as well as cardiac substructure segmentation [97]. Other significant developments include MT-UDA [98], which introduces a multi-teacher framework, and [99] further integrates frequency and spatial domain through multi-teacher distillation. Moreover, [100] explores a ’student-to-partner’ paradigm during various training stages.
@@
Graph-based Methods. Graph-based methods are increasingly utilized in cross-domain medical image analysis due to their capability to capture complex spatial structures and relationships. This approach models image elements – ranging from individual pixels to entire regions – as nodes in a graph, with edges formed based on criteria like spatial proximity and similarity in intensity or texture. The core of this method involves a graph $G=(V,E)$, with $V$ representing the vertices and $E$ the edges, which are weighted according to the mentioned criteria. This setup facilitates the use of graph convolutional networks (GCNs) [101, 102, 103], which leverage the graph structure for learning, described mathematically as:
$H^{(l+1)}=\sigma\left(D^{-\frac{1}{2}}AD^{-\frac{1}{2}}H^{(l)}W^{(l)}\right)$ (8)
@@
where $H^{(l)}$ represents the features at each node in layer $l$, $A$, $D$ is adjacency and degree matrix, $W^{(l)}$ is the weight matrix for layer $l$, $\sigma$ denotes the activation function. This process effectively leverages node features and graph topology for a comprehensive analysis. Applications of this method include feature disentanglement [104] for domain-invariant learning [105] with GCN, graph Laplacian decomposition for brain imaging alignment across domains [106], attention-guided GCN for identifying major depressive disorder [107], and a class-aware GCN classifier with domain-specific features for predicting lymph node metastasis in gastric cancer [108]. Other notable implementations like [109] extends beyond traditional methods by incorporating an online sub-graph scheme, [110] employs GCNs with a meta-learning strategy targeting at small-sized pancreatic cancer features. Studies like [111, 112] focus on enhancing feature alignment and understanding inter-category relationships using graph-based techniques.
IV-C3 Optimization Strategy
@@
Statistical Discrepancies Minimization. Quantifying and subsequently minimizing the statistical discrepancies between source and target domain feature spaces serves as a key approach. This paradigm, rooted in the hypothesis that reducing such discrepancies aids in model adaptation, predominantly employs measures like Kullback-Leibler (KL) divergence, Maximum Mean Discrepancy (MMD), and Correlation Alignment (CORAL). For example, the adaptation with MMD [113] between source and target domains can be mathematically formulated as:
$\small\mathcal{L}(D_{s},D_{t})$ $=\frac{1}{n_{s}^{2}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{s}}k(z_{s}^{i},z_{s}^{j})+\frac{1}{n_{t}^{2}}\sum_{i=1}^{n_{t}}\sum_{j=1}^{n_{t}}k(z_{t}^{i},z_{t}^{j})$
$-\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}k(z_{s}^{i},z_{t}^{j})$ (9)
@@
In the medical adaptation field, [114] explores the utility of MMD for domain adaptation in breast and thyroid lesions in ultrasound images.[115] leverages KL divergence to synchronize the prior distribution of the synthesized and the real target distribution. [116] estimates the mutual information with KL divergence between the reconstruction output and segmentation result, so as to benefit each other. [117] enforces recursively conditional Gaussian (RCG) as the joint distribution prior, inheriting the closed form of the KL divergence term in the variational objective to make large-sacle tasks computationally tractable. [118] uses MultiKernel Maximum Mean Discrepancy (MK-MMD) in aligning feature distributions in breast ultrasound images. Beyond these widely adopted metrics, novel metrics have been developed to suit specific medical tasks. [119] proposes the Characteristic Function (CF) Distance, transforming feature distributions to frequency domain for discrepancy calculations. [120] introduces Domain Sanity Loss, focusing on anatomical features like centroid distance and plausibility in vertebrae prediction.
V Federated Learning
Federated Learning is a pivotal model training approach designed to handle data heterogeneity while preserving the privacy of each client. It is particularly valuable in MedIA for alleviating data distribution shifts, allowing for collaborative enhancements across multiple healthcare institutions without the need to centralize their data. This decentralized method ensures the privacy of patient information, making it a practical solution for scenarios where medical data cannot be openly shared. One prominent example of Federated Learning in practice is FedAvg [25], which forms the basis for many modern implementations. In this model, each participating institution trains a local model on its own data, thereby maintaining the confidentiality of sensitive information. These institutions then send their model updates – commonly in the form of weights or gradients – to a central server. The server aggregates these updates to enhance the global model, which is shared back with all participating institutions after a few iterations. The mathematical formulation is detailed as follows:

### v2

@@ 5.2. Semi-supervised Joint Training
5.2.1. Data Management
5.2.1.1 Pseudo-labeling.
Pseudo-labeling is a powerful technique in semi-supervised joint training that leverages large volumes of unlabeled data to enhance model training. This method involves generating artificial labels for unlabeled data based on the most confident predictions of the model, thereby expanding the training dataset effectively. In domain adaptation, its effectiveness hinges on prioritizing model’s predictions exceeding a defined threshold that are more likely to be accurate:
(1) $\hat{y}_{u}=\begin{cases}\arg\max q_{\theta}(x_{u}),&\text{if }\max q_{\theta}(x_{u})>\tau\\ \text{ignore},&\text{otherwise}\end{cases}$
where $x_{u}$ represents an unlabeled sample from the target domain, $q_{\theta}(x_{u})$ denotes the model’s probabilistic predictions, $\hat{y}_{u}$ is the pseudo-label assigned to $x_{u}$, and $\tau$ is a threshold defining the confidence level above which the labels are considered reliable. The loss is computed as:
(2) $\mathcal{L}_{u}=\sum_{x_{u}\in\mathcal{X}_{u}}\mathbb{I}\left(\max q_{\theta}(x_{u})>\tau\right)\cdot\mathcal{L}(\hat{y}_{u},q_{\theta}(x_{u}))$
@@ 5.2.1. Data Management
(3) $\mathcal{L}_{total}=\mathcal{L}_{l}(\mathcal{Y}_{l},q_{\theta}(\mathcal{X}_{l}))+\beta\mathcal{L}_{u}(\mathcal{\hat{Y}}_{u},q_{\theta}(\mathcal{X}_{u}))$
@@
To mitigate the risk of error propagation, which can occur if incorrect labels are used for training, enhancements are made to ensure the quality of these pseudo-labels. For example, (Ghamsarian et al., 2023) employs transformation-invariant, highly-confident predictions in the target dataset for self-training purposes, ensuring that the model is less likely to learn from noisy, less reliable labels. Meanwhile, (Yao et al., 2022a) enhances the robustness of pseudo-labeling by calculating the variance between the original image and its Fourier-transformed counterpart, providing a more stable basis for generating reliable pseudo-labels. These strategies significantly improve the utility of pseudo-labeling, making it a vital tool for utilizing unlabeled data in domain adaptation.
5.2.2. Model Design
5.2.2.1 Self-ensembling
. Self-ensembling is an advanced learning strategy that effectively exploits both labeled and unlabeled data with the consistency between models. This method trains multiple versions of a model, each subjected to distinct input perturbations, and employs consistency regularization to ensure uniform predictions across these variations. Typically, this technique under SSDA setting is implemented via a “teacher-student” model, where a stable, pre-trained “teacher” model $q_{\theta}^{tc}$ guides a less-trained “student” model $q_{\theta}^{st}$. The overall training process is governed by two primary loss functions:
$\mathcal{L}_{{total}}=\mathcal{L}_{l}(\mathcal{Y}_{l},q_{\theta}^{tc}(\mathcal{X}_{l}))+\lambda\mathcal{L}_{{cons}}(q_{\theta}^{tc}(\mathcal{X}_{u})),q_{\theta}^{st}(\mathcal{X}_{u}^{aug}))$
@@ 5.2.2.1 Self-ensembling
Following this framework, (Li et al., 2020e, d) introduce this techniques into achieveing semi-supervised domain adaptation. The core innovation of these frameworks lies in the strategic employment of dual-teacher models: one teacher model enhances intra-domain knowledge through self-ensembling techniques, while the other facilitates inter-domain knowledge transfer using image translation models such as CycleGAN (Zhu et al., 2017). This approach leverages the consistency of model outputs across different views of the same data, enhancing the model’s ability to generalize across diverse scenarios.
5.2.2.2 Adversarial Learning.
Adversarial leaning is fundamentally embodied by Generative Adversarial Networks (GANs) (AlAmir and AlGhamdi, 2022). It can be conceptualized as a game between two players:
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$
where generator (G) aims to produce data that is indistinguishable from real data by transforming input noise $z$, sampled from a noise distribution $p_{z}(z)$, and discriminator (D) aims to correctly classify real data $x$ and generated data $G(z)$. Real data $x$ is sampled from the true data distribution $p_{\text{data}}(x)$. This interaction forms a min-max game where the generator seeks to deceive the discriminator into accepting its outputs as real, while the discriminator improves at identifying the differences between real and generated data. Through this adversarial process, $G$ refines its outputs to reduce discrepancies, indirectly generating domain-invariant features for domain adaptation. As one notable method, COVID-DA (Zhang et al., 2020a) is designed to distinguish between closely related conditions such as pneumonia and COVID-19, particularly when labeled data is scarce. This method uses a unique classifier separation scheme along with an adversarial network to overcome the task difference and domain discrepancy simultaneously.
5.2.2.3 Novel Training Strategies.
(Liu et al., 2022e) explores the richness of multi-modal data through a novel asymmetric co-training approach. By segmenting the learning process into two distinct components that each addresses specific aspects of domain adaptation and semi-supervised learning task, this strategy avoids the source domination thus facilitates more effective domain adaptation.
5.2.3. Optimization Strategy
5.2.3.1 Metric Learning.
Metric Learning within the context of semi-supervised joint training is distinctly innovative. In this scenario, (Basak and Yin, 2023) adopts a metric learning strategy characterized by a disentangled paradigm. This approach separates style and content into distinct embedding spaces. Such separation facilitates independent contrastive learning for each aspect, allowing the model to adapt more effectively to variations in data distributions.
5.3. Unsupervised Joint Training
@@
5.3.1. Data Management
5.3.1.1 Cross-Modal Translation.
Cross-modal translation, employing techniques such as Generative Adversarial Networks (GANs) and frequency-based methods, is pivotal in transforming how we address domain differences by converting data from the source domain $S$ to closely resemble the target domain $T$. This transformation is formalized as:
(4) $\mathcal{\tilde{X}}=G(\mathcal{X}_{s};\theta_{G})\quad\text{where}\quad G:\mathcal{D}_{s}\rightarrow\mathcal{D}_{t}$
@@ 5.3.1. Data Management
(5) $\min_{\theta_{G}}d(\mathcal{\tilde{X}},\mathcal{X}_{t})=\min_{\theta_{G}}d(G(\mathcal{X}_{s};\theta_{G}),\mathcal{X}_{t})$
@@
The application of GANs has evolved significantly, beginning with methods like the pixel-to-pixel (pix2pix) translation and advancing to more complex implementations. For instance, (Li et al., 2021b) utilizes an end-to-end unsupervised method for enhancing contrast in cataract fundus image based on pix2pix framework. Then, MADGAN (Han et al., 2021) breaks the constraints of paired images, contributing to anomaly detection in complex brain structures. SASAN (Tomar et al., 2021) takes a further step that incorporating self-attention modules in its GANs, enhancing focus on specific anatomical details during image translation. Subsequently, the utilization of CycleGANs (Zhu et al., 2017) marks another significant advancement, enabling unpaired image translations for cross-domain chest X-ray disease recognition (Tang et al., 2019; Sanchez et al., 2022) and hip joint bone segmentation (Zeng et al., 2021). This translation process is further refined in (Zou et al., 2020) with dual-scheme (source-target/target-source) fusion and (Dong et al., 2020) with attention mechanism. Integrating disentangled representations into GAN frameworks, as seen in (Chen and Wang, 2021; Peng et al., 2022; Wang and Zheng, 2022; Yang et al., 2019a; Jiang and Veeraraghavan, 2020; Yao et al., 2022b), significantly advances domain adaptation by separating content from style, enhancing adaptation efficiency. Complementing the adversarial nature of GANs, frequency-based methods (Zhang et al., 2022a; Hu et al., 2022b) introduce a novel perspective. They assume that the style information is stored in low frequency components and high frequency components represents more structural information, and thus translate the images by replacing the low frequency components. Finally, techniques like singular value decomposition for noise adaptation in retinal OCT images (Koch et al., 2022) highlight the innovation and adaptability in this field, which is tailored to specific imaging modalities or diagnostic requirements.
5.3.1.2 Pseudo-labeling.
@@
In UDA, as all target data labels are unknown, it becomes more challenging to make accurate pseudo-label predictions using traditional techniques. Research has since advanced the pseudo-labeling concept by integrating pseudo-labeling and adversarial learning to enhance the process (Xie et al., 2022). Subsequent studies have built on this foundation, each offering unique improvements to address issues such as noisy labels (Du and Liu, 2021) and enhancing label reliability through methods like iterative self-training (Shin et al., 2022), contrastive learning (Liu et al., 2022f), and entropy constraints (Feng et al., 2023a). The specialized applications of pseudo-labeling are further explored in studies (Li et al., 2022c; Cho et al., 2022; Liu et al., 2021f). For example, (Li et al., 2022c) focuses on nuclei instance segmentation and classification, utilizing pseudo-labels derived from prototype features. (Cho et al., 2022) breaks new ground in cell detection with a pseudo-cell-position heatmaps. (Liu et al., 2021f) innovates by incorporating pseudo-labeling into tagged-to-cine MRI synthesis task, employing a Bayesian uncertainty mask for selective pseudo-label generation.
5.3.2. Model Design
5.3.2.1 Adversarial Learning.
@@ 5.3.2. Model Design
Adversarial Learning is widely used for the implicit alignment between domains at feature or/and pixel level due to the absence of target labels. At the feature level, techniques such as the plug-and-play adversarial domain adaptation network (PnP-AdaNet) (Dou et al., 2019b) aligns features across different scales for segmentation tasks. Similarly, (Jiang et al., 2022a) aligns extracted contents for cross-modality segmentation. Other studies focus on prediction space alignment at the pixel level for various medical imaging tasks (Liu et al., 2022c; Lei et al., 2021; Yan et al., 2019; Shen et al., 2020). Integrated approaches that apply adversarial training at both feature and output levels are explored in studies like (Chen et al., 2019a, 2020; Shin et al., 2021; Liu et al., 2022g; Su et al., 2023b; Calisto and Lai-Yuen, 2022; Bian et al., 2022). Innovations in this field also include enhanced discriminators and local discriminators that focus on specific region alignment, introducing spatial-aware and class-specific attentions to refine the adversarial loss and improve model adaptability across domains (Pham et al., 2021; Yoo et al., 2021; Liu et al., 2021d; Chen et al., 2022a).
5.3.2.2 Self-ensembling.
@@
Initial studies by (Perone et al., 2019) applied self-ensembling to gray matter MRI segmentation. Subsequent applications include breast MRI segmentation (Kuang et al., 2023) and pose estimation in operating rooms (Srivastav et al., 2022). More advanced techniques combine adversarial training and self-ensembling for addressing domain shifts in cross-institutional gliomas studies (Shanis et al., 2019), optic disc and cup segmentation (Liu et al., 2019) as well as cardiac substructure segmentation (Zhao et al., 2022). Other significant developments include MT-UDA (Zhao et al., 2021), which introduces a multi-teacher framework, and (Liu et al., 2023d) further integrates frequency and spatial domain through multi-teacher distillation. Moreover, (Hong et al., 2022a) explores a ’student-to-partner’ paradigm during various training stages.
5.3.2.3 Graph-based Methods.
@@
Graph-based methods are increasingly utilized in cross-domain medical image analysis due to their capability to capture complex spatial structures and relationships. This approach models image elements – ranging from individual pixels to entire regions – as nodes in a graph, with edges formed based on criteria like spatial proximity and similarity in intensity or texture. The core of this method involves a graph $G=(V,E)$, with $V$ representing the vertices and $E$ the edges, which are weighted according to the mentioned criteria. This setup facilitates the use of graph convolutional networks (GCNs) (Kipf and Welling, 2016; Veličković et al., 2017; Guo et al., 2024), which leverage the graph structure for learning, described mathematically as:
(6) $H^{(l+1)}=\sigma\left(D^{-\frac{1}{2}}AD^{-\frac{1}{2}}H^{(l)}W^{(l)}\right)$
@@
where $H^{(l)}$ represents the features at each node in layer $l$, $A$, $D$ is adjacency and degree matrix, $W^{(l)}$ is the weight matrix for layer $l$, $\sigma$ denotes the activation function. This process effectively leverages node features and graph topology for a comprehensive analysis. Applications of this method include feature disentanglement (Guo et al., 2022) for domain-invariant learning (Mahapatra et al., 2022) with GCN, graph Laplacian decomposition for brain imaging alignment across domains (Gopinath et al., 2020), attention-guided GCN for identifying major depressive disorder (Fang et al., 2023), and a class-aware GCN classifier with domain-specific features for predicting lymph node metastasis in gastric cancer (Zhang et al., 2022c). Other notable implementations like (Liu et al., 2021e) extends beyond traditional methods by incorporating an online sub-graph scheme, (Li et al., 2021a) employs GCNs with a meta-learning strategy targeting at small-sized pancreatic cancer features. Studies like (Liu et al., 2021c, b) focus on enhancing feature alignment and understanding inter-category relationships using graph-based techniques.
5.3.3. Optimization Strategy
5.3.3.1 Statistical Discrepancies Minimization.
@@ 5.3.3. Optimization Strategy
Quantifying and subsequently minimizing the statistical discrepancies between source and target domain feature spaces serves as a key approach. This paradigm, rooted in the hypothesis that reducing such discrepancies aids in model adaptation, predominantly employs measures like Kullback-Leibler (KL) divergence, Maximum Mean Discrepancy (MMD), and Correlation Alignment (CORAL). For example, the adaptation with MMD (Long et al., 2016) between source and target domains can be mathematically formulated as:
$\small\mathcal{L}(D_{s},D_{t})=\frac{1}{n_{s}^{2}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{s}}k(z_{s}^{i},z_{s}^{j})+\frac{1}{n_{t}^{2}}\sum_{i=1}^{n_{t}}\sum_{j=1}^{n_{t}}k(z_{t}^{i},z_{t}^{j})-\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}k(z_{s}^{i},z_{t}^{j})$
@@
In the medical adaptation field, (Gao et al., 2022) explores the utility of MMD for domain adaptation in breast and thyroid lesions in ultrasound images.(Hu et al., 2022a) leverages KL divergence to synchronize the prior distribution of the synthesized and the real target distribution. (Lu et al., 2022) estimates the mutual information with KL divergence between the reconstruction output and segmentation result, so as to benefit each other. (Liu et al., 2022b) enforces recursively conditional Gaussian (RCG) as the joint distribution prior, inheriting the closed form of the KL divergence term in the variational objective to make large-sacle tasks computationally tractable. (Xu et al., 2022b) uses MultiKernel Maximum Mean Discrepancy (MK-MMD) in aligning feature distributions in breast ultrasound images. Beyond these widely adopted metrics, novel metrics have been developed to suit specific medical tasks. (Wu and Zhuang, 2020) proposes the Characteristic Function (CF) Distance, transforming feature distributions to frequency domain for discrepancy calculations. (Sager et al., 2022) introduces Domain Sanity Loss, focusing on anatomical features like centroid distance and plausibility in vertebrae prediction.
Figure 4. Federated Learning in MedIA. It facilitates collaborative model training across healthcare institutions while preserving data privacy. By maintaining data decentralization, it effectively addresses data distribution shifts without centralizing sensitive medical information, ensuring privacy and enhancing model robustness.
6. Federated Learning
Federated Learning is a pivotal model training approach designed to handle data heterogeneity while preserving the privacy of each client. It is particularly valuable in MedIA for alleviating data distribution shifts, allowing for collaborative enhancements across multiple healthcare institutions without the need to centralize their data (see Fig. 4). This decentralized method ensures the privacy of patient information, making it a practical solution for scenarios where medical data cannot be openly shared. One prominent example of Federated Learning in practice is FedAvg (McMahan et al., 2017), which forms the basis for many modern implementations. In this model, each participating institution trains a local model on its own data, thereby maintaining the confidentiality of sensitive information. These institutions then send their model updates – commonly in the form of weights or gradients – to a central server. The server aggregates these updates to enhance the global model, which is shared back with all participating institutions after a few iterations. The mathematical formulation is detailed as follows: